Bramble Corp. erected and placed into service an offshore oil platform on January 1, 2020, at a cost of $8 million. Bramble is legally required to dismantle and remove the platform at the end of its 7-year useful life. Bramble estimates that it will cost $1 million to dismantle and remove the platform at the end of its useful life and that the discount rate to use should be 6%. Use (a) factor Table A.2, (b) a financial calculator, or (c) Excel function PV. Ignore production related costs for this question.

Prepare any necessary adjusting entries that are associated with the asset retirement obligation and related expenses at December 31, 2020, assuming that Bramble follows IFRS. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 5,275. Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Depreciation Expense  
Accumulated Depreciation  
(To record depreciation expense)
Interest Expense  
Asset Retirement Obligation  
(To record interest expense)

Prepare any necessary adjusting entries that are associated with the asset retirement obligation and related expenses at December 31, 2020, assuming that Bramble follows ASPE. (Round factor values to 5 decimal places, e.g. 1.25124 and final answer to 0 decimal places, e.g. 5,275. Credit account titles are automatically indented when the amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Depreciation Expense  
Accumulated Depreciation  
(To record depreciation expense)
Accretion Expense  
Asset Retirement Obligation  
(To record interest expense)
Answer #1

IFRS and ASPE

Cost = $8,000,000

Present value of dismantling costs = $1,000,000 * 0.665 = $665,057.1136

Total cost = $8,665,057.1136

Depreciation per year = $8,665,057.1136 / 7 years = $1,237,865

Dr Depreciation expense $1,237,865

Cr Accumulated depreciation $1,237,865

Interest expense = $665057.1135 * 6% = $39903

Dr Interest expense $39903

Cr Asset retirement obligation $39903